# Quantum Computing Education in Latin America: Experiences and Strategies

### v1

@@ Abstract
Quantum computing is a rapidly advancing field facing a significant shortage of qualified experts, creating an urgent demand for skilled professionals. In Latin America, quantum education remains in its early stages, further exacerbating the regional talent gap due to deficiencies in educational infrastructure and limited financial support. This work presents the efforts of the Quantum Computing and Artificial Intelligence research group at Universidad Nacional de Colombia to integrate quantum computing into higher education using the European Competence Framework for Quantum Technologies, which offers a standardized approach to defining competency requirements and assessing essential skills in the field. We propose introductory courses aligned with the framework’s guidelines, designed to meet industry standards and reach a broad audience. Additionally, we introduce supplementary resources, including team dynamics and evaluation methodologies, to enhance the educational ecosystem. These initiatives aim to create a sustainable and comprehensive quantum education model across Latin America, to close the regional skills gap, foster inclusivity, and prepare a diverse workforce to contribute innovative solutions on the global stage.
@@
Quantum Computing, Education, Latin America, Workforce Development, Competence Framework.
@@ I Introduction
Quantum computing (QC), an emerging technology that is rapidly advancing in the research field [1], holds the promise to revolutionize society as a significant breakthrough in modern IT in the coming years [2], [3]. Its importance is evident from the substantial investments by developed nations and leading companies worldwide in research and development, aimed at achieving practical quantum computers. This captures the attention of individuals, countries, and governments alike due to its promising applications in almost every area imaginable [2], [4]. This field is profoundly multidisciplinary, requiring expertise in various fields for its development. As quantum technologies (QT) become increasingly important in industry, there is a growing demand for a skilled quantum workforce [5]. However, there is a significant global shortage of qualified professionals in QC, creating an urgent need for a new cohort of engineers, scientists, and programmers who can leverage quantum computers to tackle real-world challenges and bring innovative problem-solving skills to the field [6].
@@
While the workforce development issue is global, Latin America lags behind due to deficiencies in its educational systems and lack of financial support, with knowledge primarily stemming from individual efforts. On one hand, QC is increasingly gaining recognition within Latin America, and the region is following the lead of other countries through various events such as conferences and seminars. Additionally, there are organizations and communities that provide a supportive environment and guidance on the topic. But on the other hand, the situation in higher education is less favorable. Although the community works to raise awareness and provide basic tools to build expertise, university-level courses are essential for individuals to be qualified and competitive on a global scale. While there are courses and master’s degrees at various universities that cover quantum-related subjects, they are neither common nor widely explored, focused specifically on QC, or accessible to a broader audience. Furthermore, the lack of governmental support results in fewer companies and institutions providing a rigorous environment for research and practical application, making it challenging for students to apply theoretical knowledge in real-world scenarios. According to the 2023 Global Quantum Computing Map [8], the United States, China, and Europe are leading the development of QC, while Japan, South Korea, Singapore, and Australia are also active participants in the global race. Venegas-Gomez [9] presents a global effort map for public funding in QT, an effort that has been enhanced by the establishment of national programs worldwide. However, no Latin American country shows significant investments in QC that would warrant notable mention on the map, nor are they included in the global distribution of companies working in QT per country.
@@ I-A The European Competence Framework for Quantum Technologies
The data underscores not only the urgency of preparing the future workforce, but also the need of tailored educational programs addressing emerging demands [10]. It is crucial to understand the specific qualification required by the industry for various job roles to create targeted courses, specialized tracks, and degree programs at universities which enhance students’ competitiveness for quantum-related positions [11]. To achieve this, the European Competence Framework for Quantum Technologies (ECFQT) [12] was developed following a study aimed at collecting and identifying the competences, knowledge, and skills relevant for the emerging quantum workforce and its development. This framework establishes standardized competency requirements for diverse roles in QT, ensuring transparency, scalability, and comparability in defining essential skills, enabling consistent assessment across different courses or qualification [13]. Therefore, one viable solution for Latin American countries is to leverage this framework as a guide to ensure a comprehensive and standardized approach to QC education.
Addressing this issue, a research group is dedicated to making QC education open and available to all students in Latin America.
@@ I-B UNAL’S Quantum Computing Research Group
The Quantum Computing and Artificial Intelligence Research Group at the Universidad Nacional de Colombia (UNAL) is at the forefront of quantum computing research in Latin America. As one of the largest research groups in the country, with over 100 students at various educational levels, they are involved in cutting-edge research projects, publications, conferences, and collaborations with industry, working to integrate QC in the region. The group meets regularly to study key research papers, deliver seminars, and host talks covering a wide range of quantum topics, while maintaining strong connections with researchers and institutions in Europe and the United States.
What sets this group apart is its commitment to open science and education. As an inclusive community, it welcomes individuals regardless of academic background and is particularly dedicated to making quantum education accessible across Latin America, where resources in Spanish are scarce or often come with a price barrier, as most educational materials—ranging from formal courses, textbooks, and self-paced online courses, to video series, programming tutorials, workshops, and games— are predominantly restricted to the English language [14], further limiting access for non-English speakers in the region. They consistently offer free courses via Twitch and YouTube, which now have thousands of views and subscribers, making high-quality information available. Their course offerings include “Quantum Computing,” “Artificial Intelligence: Exploring Neural Networks,” and “From Qubits to Qudits.” The group is actively engaging in more national and international collaborations, aiming to build a robust quantum ecosystem in Colombia and Latin America. They are planning to expand their course offerings in the future and intend to use the ECFQT to ensure that the education provided aligns with global needs.
@@ II Methodology
We will now present the implementation of the European Competence Framework for Quantum Technologies to develop two initial quantum computing courses. To begin, we will first explore its structure.
The framework consists of three main parts. First, a content map (Fig. 1) with 8 domains and 42 subdomains, which provides a graphical overview of the structure and topics of QT.
Figure 1: Framework’s domains overview with topics. Source: European Competence Framework for Quantum Technologies.
The content map also contains proficiency levels from A1 to C2 (see Table I), analogously to the CERF levels for languages, to tailor education to specific target groups’ needs and to determine job profiles. The depth of knowledge required for each level are described, along with the time it takes to achieve it. These levels apply to three proficiency areas that encompass different competences, each being color-coded according to the domains it covers:
I. Quantum Concepts
II. QT Hardware and Software Engineering
III. QT Applications and Strategies
Thus, the first area (in green) is associated with domains 1 and 2, the second area (in blue) corresponds to domains 3, 4, 5, 6, and 7, and the third area (in red) is correlated with domain 8.
@@ III. QT Applications and Strategies
Proficiency Level Description
@@
Using the proficiency levels and areas, the proficiency triangles (Part 2) illustrate the general qualifications covered by a person or addressed in a course, as shown in Fig. 2. This figure illustrates a qualification profile created using the Qualification Profile Creation Tool, provided by the ECFQT authors, where the individual achieved C1, B2, and C2 proficiency levels in the three respective areas.
@@
Finally, Part 3 consists of qualification profiles (from P1 to P9) that provide an overview of common personal QT-specific qualifications relevant to industry. The framework shows additional insights on the profiles, including descriptions of the highest levels addressed, a general explanation of the profile, example personas and related recommendations.
@@ II-A Proposed Courses and Implementation Strategy
1. The proficiency triangle for the course’s specific qualification profile, illustrating the proficiency level students are expected to reach in the three main areas by the end of each course.
2. A general overview of the expected outcomes, outlining what students should be able to accomplish with the proficiency levels attained in each of the three areas covered by the framework.
3. A selection of topics and subtopics from the 42 available options, carefully chosen to align with the course objectives.
4. Following the framework’s guidelines to combine proficiency levels with an appropriate selection and adaptation of content (sub)domains to define a course’s objectives or an individual’s qualifications, we developed a detailed target profile. This profile outlines the specific knowledge and skills students will acquire in each area. The qualification profile was designed based on the framework’s recommendations, while being tailored to fit the objectives of each course.
@@
    • Duration: Four weeks (one hour per week).
@@
    Description
    This introductory course provides students with an initial exposure to the current state of QT. It offers a broad perspective on how different quantum technologies interconnect (not solely covering QC), and is suitable for non-STEM degrees and professionals from diverse fields working in quantum-related projects, providing basic knowledge to those without a technical background.
    Objectives
@@
    Adaptation to the ECFQT
    For this course, the combination of the P1 qualification profile with the Quantum Background domain (1) from the content map leads to the target profile of General QT aware person. The description and qualifications are shown in Table II. Fig. 3 illustrates the proficiency triangle and its levels description, and Fig. 4 shows the chosen focus contents for the course.
    Figure 3: Target proficiency triangle for the first course and general descriptions of the target proficiency level for each area. Source: European Competence Framework for Quantum Technologies.
    Figure 4: Topics and subtopics to be covered in the first course. Adapted from the European Competence Framework for Quantum Technologies.
    TABLE II: Persona’s specific description and specific qualifications (First Course)
@@
    • Duration: Sixteen hours (4 hours per week).
@@
    Description
    This course shifts the focus more towards QC, offering deeper insights into its applications and challenges. The course combines both theoretical and practical aspects of QC, providing opportunities to interact with professionals and engage in problem-solving tasks. Despite the advanced content and mathematical formalism, the course is designed to be relatively accessible as a basic course. Combines content from the ECFQT and a selected guidebook to ensure a QC-focused curriculum.
    Objectives
@@
    Adaptation to the ECFQT
    For this course, the combination of the P3 qualification profile with the Quantum Computing and Simulations domain (5) from the content map leads to the target profile of QC literate person. The description and qualifications are shown in Table III. Fig. 5 illustrates the proficiency triangle and its levels description, and Fig. 6 shows the chosen focus contents.
    Figure 5: Target proficiency triangle for the second course and general descriptions of the target proficiency level for each area. Source: European Competence Framework for Quantum Technologies.
    Figure 6: Topics and subtopics to be covered in the second course. Adapted from the European Competence Framework for Quantum Technologies.
@@ II-B Additional Proposals
Work Teams
Analogously to the qualification profiles and example personas given in Part 3 of the ECFQT, examples of real work teams could be integrated into the courses to illustrate applicability and provide students with real-life scenarios of QC applications and understand the tangibility of their future work. A starting point is to use the already existing example personas in the framework, built by combining qualification profiles with a selection from the content map to lead to titles that specify the focus, and later new profiles or professions outside the field can be included. Next, we present a work team example, which does not necessarily reflect a real-life situation but serves for illustrative purposes, to demonstrate how the framework’s concepts could be applied to profile team members.
@@
I. Application’s Description: The project focuses on leveraging QC techniques to optimize financial portfolios in real-time. Traditional portfolio optimization methods face limitations in handling large datasets and complex optimization criteria.
II. Objective: Develop and implement quantum algorithms that can optimize financial portfolios more effectively than classical methods, to maximize returns while minimizing risks, considering various financial constraints and market dynamics.
III. Team Collaboration
@@
IV. Workflow Overview
Requirements Analysis: The Financial Market Analyst collaborates with stakeholders to define optimization objectives and constraints.
$\downarrow$
Algorithm Development: The Quantum Algorithm Computer Scientist designs and implements quantum algorithms based on defined optimization criteria.
$\downarrow$
Integration and Testing: The Quantum Computing Specialist ensures seamless integration of algorithms with quantum hardware and conducts performance testing.
$\downarrow$
Market Assessment: The Business Strategist evaluates market opportunities, develops business plans, and identifies potential partners or clients.
Evaluation
Based on the framework’s proficiency level descriptions for the three areas - Quantum Concepts, QT HW and SW Engineering, and QT Applications and Strategies - which detail the expected knowledge and skills, course evaluation tasks can be developed to assess student learning. By integrating these tasks, clear metrics and assessment tools can be established to measure both the knowledge (K) and skills (S) outlined in the framework, ensuring a thorough assessment of student performance.
@@ I. Quantum Concepts (Target level: A1 - Awareness)
  • Example: “What is the concept of quantum superposition? (a) The ability of a particle to be in multiple states at once (b) The ability of a photon to interact in such a way that its quantum state cannot be described independently.”

### v2

@@ Abstract
Quantum computing is a rapidly advancing field facing a significant shortage of qualified experts, creating an urgent demand for skilled professionals. In Latin America, quantum education remains in its early stages, further exacerbating the regional talent gap due to deficiencies in educational infrastructure and limited financial support. This work presents the efforts of the Quantum Computing Research Group at Universidad Nacional de Colombia to integrate quantum computing into higher education using the European Competence Framework for Quantum Technologies, which offers a standardized approach to defining competency requirements and assessing essential skills in the field. We propose introductory courses aligned with the frameworks guidelines, designed to meet industry standards and reach a broad audience. Additionally, we introduce supplementary resources, including team dynamics and evaluation methodologies, to enhance the educational ecosystem. These initiatives aim to create a sustainable and comprehensive quantum education model across Latin America, to close the regional skills gap, foster inclusivity, and prepare a diverse workforce to contribute innovative solutions on the global stage.
@@
Quantum Education, Latin America, Workforce Development, Competence Framework.
@@ I Introduction
Quantum computing (QC), an emerging technology that is rapidly advancing in the research field [1], holds the promise to revolutionize society as a significant breakthrough in modern IT in the coming years [2], [3]. Its importance is evident from the substantial investments by developed nations and leading companies worldwide in research and development, aimed at achieving practical quantum computers. This captures the attention of individuals, countries, and governments alike due to its promising applications in almost every area imaginable [2], [4].
This field is profoundly multidisciplinary, requiring expertise in various fields for its development. As quantum technologies (QT) become increasingly important in industry, there is a growing demand for a skilled quantum workforce [5]. However, there is a significant global shortage of qualified professionals in QC, creating an urgent need for a new cohort of engineers, scientists, and programmers who can leverage quantum computers to tackle real-world challenges and bring innovative problem-solving skills to the field [6].
@@
While the workforce development issue is global, Latin America lags behind due to deficiencies in its educational systems and lack of financial support, with knowledge primarily stemming from individual efforts. On one hand, QC is increasingly gaining recognition within Latin America, and the region is following the lead of other countries through various events such as conferences and seminars. Additionally, there are organizations and communities that provide a supportive environment and guidance on the topic. But on the other hand, the situation in higher education is less favorable.
Although the community works to raise awareness and provide basic tools to build expertise, university-level courses are essential for individuals to be qualified and competitive on a global scale. While there are courses and master’s degrees at various universities that cover quantum-related subjects, they are neither common nor widely explored, focused specifically on QC, or accessible to a broader audience. Furthermore, the lack of governmental support results in fewer companies and institutions providing a rigorous environment for research and practical application, making it challenging for students to apply theoretical knowledge in real-world scenarios.
According to the $20232023$ Global Quantum Computing Map [8], the United States, China, and Europe are leading the development of QC, while Japan, South Korea, Singapore, and Australia are also active participants in the global race. Venegas-Gomez [9] presents a global effort map for public funding in QT, an effort that has been enhanced by the establishment of national programs worldwide. However, no Latin American country shows significant investments in QC that would warrant notable mention on the map, nor are they included in the global distribution of companies working in QT per country.
@@ I-A The European Competence Framework for Quantum Technologies
The data underscores not only the urgency of preparing the future workforce, but also the need of tailored educational programs addressing emerging demands [10]. It is crucial to understand the specific qualification required by the industry for various job roles to create targeted courses, specialized tracks, and degree programs at universities which enhance students’ competitiveness for quantum-related positions [11]. To achieve this, the European Competence Framework for Quantum Technologies (CFQT) [12] was developed following a study aimed at collecting and identifying the competences, knowledge, and skills relevant for the emerging quantum workforce and its development [5, 10]. This framework establishes standardized competency requirements for diverse roles in QT, ensuring transparency, scalability, and comparability in defining essential skills, enabling consistent assessment across different courses or qualification [13, 14]. Therefore, one viable solution for Latin American countries is to leverage this framework as a guide to ensure a comprehensive and standardized approach to QC education. Addressing this issue, a research group is dedicated to making QC education open and available to all students in Latin America.
@@ I-B UNAL’S Quantum Computing Research Group
The Quantum Computing Research Group at the Universidad Nacional de Colombia (UNAL) is at the forefront of quantum computing research in Latin America. As one of the largest research groups in the country, with over $7070$ students at various educational levels, they are involved in cutting-edge research projects, publications, conferences, and collaborations with industry, working to integrate QC in the region. The group meets regularly to study key research papers, deliver seminars, and host talks covering a wide range of quantum topics, while maintaining strong connections with researchers and institutions in Europe and the United States. What sets this group apart is its commitment to open science and education. As an inclusive community, it welcomes individuals regardless of academic background and is particularly dedicated to making quantum education accessible across Latin America, where resources in Spanish are scarce or often come with a price barrier, as most educational materials—ranging from formal courses, textbooks, and self-paced online courses, to video series, programming tutorials, workshops, and games— are predominantly restricted to the English language [15], further limiting access for non-English speakers in the region. They consistently offer free courses via Twitch and YouTube, which now have thousands of views and subscribers, making high-quality information available. Their course offerings include “Quantum Computing,” “Artificial Intelligence: Exploring Neural Networks,” and “From Qubits to Qudits.” The group is actively engaging in more national and international collaborations, aiming to build a robust quantum ecosystem in Colombia and Latin America. They are planning to expand their course offerings in the future and intend to use the CFQT to ensure that the education provided aligns with global needs.
@@ II Methodology
We will now present the implementation of the European Competence Framework for Quantum Technologies (version 2.5) to develop two initial quantum computing courses. To begin, we will first explore its structure. The framework consists of three main parts. First, a content map Fig. 1 with $88$ domains and $4242$ subdomains, which provides a graphical overview of the structure and topics of QT.
Figure 1: Framework domains overview with topics. Source: European Competence Framework for Quantum Technologies.
The framework also contains proficiency levels from A1 to C2 (Tab. I), analogously to the Common European Framework of Reference for Languages (CERF) [16] levels for languages, to tailor education to specific target groups’ needs and to determine job profiles. The depth of knowledge required for each level are described, along with the time it takes to achieve it. These levels apply to three proficiency areas that encompass different competences, each being color-coded according to the domains it covers:
  1. I.
    Quantum Concepts.
  2. II.
    QT Hardware and Software Engineering.
  3. III.
    QT Applications and Strategies.
Thus, the first area (green) is associated with domains $11$ and $22$, the second area (blue) corresponds to domains $33$, $44$, $55$, $66$, and $77$, and the third area (red) is correlated with domain $88$ and the application-oriented subdomains from domains $55$, $66$, and $77$.
@@
Proficiency Level Keyword
@@
Using the proficiency levels and areas, the proficiency triangles (Part $22$) illustrate the general qualifications covered by a person or addressed in a course, as shown in Fig. 2. This figure illustrates a qualification profile created using the Qualification Profile Creation Tool, provided by the CFQT authors, where the individual achieved C1, B2, and C2 proficiency levels in the three respective areas.
@@
Finally, Part $33$ consists of qualification profiles (from P1 to P9) that provide an overview of common personal QT-specific qualifications relevant to industry. The framework shows additional insights on the profiles, including descriptions of the highest levels addressed, a general explanation of the profile, example personas and related recommendations.
@@ II-A Proposed Courses and Implementation Strategy
  1. 1.
    The proficiency triangle for the course’s specific qualification profile, illustrating the proficiency level students are expected to reach in the three main areas by the end of each course.
  2. 2.
    A general overview of the expected outcomes, outlining what students should be able to accomplish with the proficiency levels attained in each of the three areas covered by the framework.
  3. 3.
    A selection of topics and subtopics from the $4242$ available options, carefully chosen to align with the course objectives.
  4. 4.
    Following the framework’s guidelines to combine proficiency levels with an appropriate selection and adaptation of content (sub)domains to define a course’s objectives or an individual’s qualifications, we developed a detailed target profile. This profile outlines the specific knowledge and skills students will acquire in each area. The qualification profile was designed based on the framework’s recommendations, while being tailored to fit the objectives of each course.
@@
    • Duration: Four weeks (two hours per week).
@@
    1. (a)
      Description: This introductory course provides students with an initial exposure to the current state of QT. It offers a broad perspective on how different quantum technologies interconnect (not solely covering QC), and is suitable for non-STEM degrees and professionals from diverse fields working in quantum-related projects, providing basic knowledge to those without a technical background.
    2. (b)
      Objectives:
@@
    3. (c)
      Adaptation to the CFQT: For this course, the primary focus aligns with the P1 qualification profile and the Quantum Background domain 1 from the content map, which defines the target profile of a General QT aware person. However, to provide a more comprehensive understanding, the course also incorporates selected subdomains related to Quantum Technology Applications (subdomains 5.6, 6.7, 7.4 and 7.6), offering participants an overview of the broader landscape of QT. Additionally, topics from domain 8 are included to highlight impact and awareness, completing a general overview. This approach ensures that learners gain foundational knowledge while appreciating the potential applications across various fields. The description and qualifications are shown in Table II. Fig. 3 illustrates the proficiency triangle and its levels description, and Fig. 4 shows the chosen focus contents for the course.
    Figure 3: Target proficiency triangle for the first course and general descriptions of the target proficiency level for each area. These concise descriptions were adapted from the more detailed descriptions in the CFQT.
    TABLE II: Persona’s description and specific qualifications (First Course)
@@
    Figure 4: Subdomains and topics to be covered in the basic course. Adapted from the European Competence Framework for Quantum Technologies.
@@
    • Duration: Sixteen weeks ($44$ hours per week).
@@
    1. (a)
      Description: This course shifts the focus more towards QC, offering deeper insights into its applications and challenges. The course combines both theoretical and practical aspects of QC, providing opportunities to interact with professionals and engage in problem-solving tasks. Despite the advanced content and mathematical formalism, the course is designed to be relatively accessible as a basic course. Combines content from the CFQT and a selected guidebook to ensure a QC-focused curriculum.
    2. (b)
      Objectives:
@@
    3. (c)
      Adaptation to the CFQT: For this course, the combination of the P3 qualification profile with the Quantum Computing and Simulations domain 5 from the content map leads to the target profile of QC literate person. The description and qualifications are shown in Table III. Fig. 5 illustrates the proficiency triangle and its levels description, and Fig. 6 shows the chosen focus contents.
    Figure 5: Target proficiency triangle for the second course and general descriptions of the target proficiency level for each area. These concise descriptions were adapted from the more detailed descriptions in the CFQT.
@@
    Figure 6: Subdomains and topics to be covered in the second course. Adapted from the European Competence Framework for Quantum Technologies.
@@ II-B Additional Proposals
  1. a.
    Work Teams: Analogously to the qualification profiles and example personas given in Part $33$ of the CFQT, examples of real work teams could be integrated into the courses to illustrate applicability and provide students with real-life scenarios of QC applications and understand the tangibility of their future work. A starting point is to use the already existing example personas in the framework, built by combining qualification profiles with a selection from the content map to lead to titles that specify the focus, and later new profiles or professions outside the field can be included. Next, we present a work team example, which does not necessarily reflect a real-life situation but serves for illustrative purposes, to demonstrate how the framework’s concepts could be applied to profile team members.
@@
    1. (a)
      Application’s Description: The project focuses on leveraging QC techniques to optimize financial portfolios in real-time. Traditional portfolio optimization methods face limitations in handling large datasets and complex optimization criteria.
    2. (b)
      Objective: Develop and implement quantum algorithms that can optimize financial portfolios more effectively than classical methods, to maximize returns while minimizing risks, considering various financial constraints and market dynamics.
    3. (c)
      Team Collaboration:
@@
  2. b.
    Workflow Overview:
    Requirements Analysis: The Financial Market Analyst collaborates with stakeholders to define optimization objectives and constraints. Algorithm Development: The Quantum Algorithm Computer Scientist designs and implements quantum algorithms based on defined optimization criteria. Integration and Testing: The Quantum Computing Specialist ensures seamless integration of algorithms with quantum hardware and conducts performance testing. Market Assessment: The Business Strategist evaluates market opportunities, develops business plans, and identifies potential partners or clients.
  3. c.
    Evaluation: Based on the framework’s proficiency level descriptions for the three areas - Quantum Concepts, QT HW and SW Engineering, and QT Applications and Strategies - which detail the expected knowledge and skills, course evaluation tasks can be developed to assess student learning. By integrating these tasks, clear metrics and assessment tools can be established to measure both the knowledge (K) and skills (S) outlined in the framework, ensuring a thorough assessment of student performance.
@@ I. Quantum Concepts (Target level: A1 - Awareness)
  • Example: “What is the concept of quantum superposition? (a) The ability of a particle to exist in a combination of different states simultaneously. (b) The ability of a photon to interact in such a way that its quantum state cannot be described independently.”